Mr.Toya had part of his $5000 savings in an account that earned 7% interest and the rest in an account that earned 9% interest. How much did he have in each account if his annual income from the total investment was $404.20?
Answer by checkley77(12844)

.09X+.07(5,000-X)=404.20
.09X+350-.07X=404.20
.02X=404.20-350
.02X=54.2
X=54.2/.02
X=$2,710 AMOUNT INVESTED @ 9%.
PROOF:
.09*2,710+.07(5,000-2,710)=404.20
243.90+.07*2,290=404.20
243.90+160.3=404.20
404.20=404.20